OpenStudy (anonymous):

Evaluate, if possible, f(x)=x^3-x ; f(x)-f(1) / x-1

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

We have \[f(x)=x^3-x\] Let's find f(1), can you find that? @schmidtdancer

OpenStudy (anonymous):

Ok yes, it would be f(1)=1-1 which is 0?

OpenStudy (ash2326):

okay, let's work to find that \[\frac{f(x)-f(1)}{x-1}\] f(1)=0 and f(x)=x^3-x \[\frac{f(x)-f(1)}{x-1}=\frac{x^3-x-0}{x-1}=\frac{x^3-x}{x-1}\] Can you simplify this?

OpenStudy (anonymous):

x^3-x/x-1?

OpenStudy (ash2326):

yeah, factor the numerator, will you?

OpenStudy (anonymous):

x(x^2-1)/x-1

OpenStudy (ash2326):

we could simplify it more \[\frac{x(x+1)(x-1)}{x-1}\] now we can cancel the (x-1) from numerator and denominator provided \(x\ne 1\) if x=1 the expression becomes \(\huge\frac 00\), it's an indeterminate form. so assuming \(x\ne 1), cancel the (x-1) s @schmidtdancer

OpenStudy (anonymous):

Yeah, ok so it is x(x+1) now

OpenStudy (ash2326):

yeah. that's our final expression under the condition that \(x\ne1\)

OpenStudy (anonymous):

Ok so I would put x(x+1) as my answer and include x doesn't equal 1?

OpenStudy (ash2326):

yeah